Find the equation of the line through (2,a) and (4,b).
I have b-a/2 =m but then i'm stuck on finding b as in y=mx+b
thank you

Answer by Alan3354(69443)

Find the equation of the line through (2,a) and (4,b).
I have b-a/2 =m but then i'm stuck on finding b as in y=mx+b
-----------
Slope m = diffy/diffx = (b-a)/2
Use y = mx + c and either point to find the y-intercept c
a = ((b-a)/2)*2 + c
a = b-a + c
c = 2a - b
---------
--> y = (b-a)x/2 + 2a - b
------------
2y = (b-a)x + 4a - 2b
(a-b)x + 2y = 4a - 2b